Tetrachloroethylene (C Cls) also known as PCE, is a solvent that was used extensively in the past to remove grease from metal machine parts. PCE wastes were often disposed in open seepage basins. Consider an industrial site in which the PCE concentration in soil below an old seepage basin is 655 mg(c)/kg(s). The PCE is degraded by naturally occurring bacteria at the site, and the process can be approximately as first order with a 1-yr half-life. 5. a. Calculate the concentration four years from now, assuming that bacterial degradation is the dominant process affecting PCE concentration in the soil b. The degradation process is actually a transformation process in which one of the C atoms in the molecule is replaced by an H atom, yielding the degradation product of C2HCls, which is another contaminant, known as TCE. What is the concentration of TCE after 5 years? Perform a mass balance for the system. c.
Answer #1

SHEET -Ba 91/2 Pard ca) 욥욧 cs = 20-48 38 : my (D/19(3)SHEET CR = 634, 5162 mg cc )N9(8) 1 66 5。2, 6438 hg (C)/lges) = Pant (C Sail Sy 8 3 83166 166 HU 36.S M 166 . Final mous 83 166 166 166 As yau can 300DEAR STUDENT, MASS BALANCE ANALYSIS CAN BE DONE IN MNY WAYS. IF YOU HAVE ANY DOUBT REGARDING MASS BALANCE ANALYSIS DONE IN PART C, FEEL FREE TO ASK. I SHALL TRY MY BEST TO RESOLVE YOUR QUERY. THANK YOU.